OpenStudy (anonymous):

HELP PLEASE!! WILL GIVE A MEDAL!! Evaluate the expression when a = –1.4 and b = -2.7 –|a| – b A.) 4.1 B.) –1.3 C.) –4.1 D.) 1.3

OpenStudy (anonymous):

@GotToLoveJenny @TechnoSoul @beautifulll

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

-4.1

OpenStudy (epoweritheta):

c

OpenStudy (anonymous):

its C.

OpenStudy (anonymous):

Absolute value of -1.4 is positive 1.4 but there is also a negative on the outside of the absolute value, and in front of it so you still get -1.4, then if you plug in -2.7 into the subtraction, a negative minus a negative is a positive so you get -1.4+2.7 and then you just add

OpenStudy (anonymous):

you'd get positive 1.3

OpenStudy (anonymous):

-(1.4)-(-2.7)=-1.4 +2.7 =1.3 as –|-1.4|=-(1.4)

OpenStudy (anonymous):

ya so you have -1.4 - - 2.7 so -1.4 +2.7= 1.3

OpenStudy (anonymous):

yes @TechnoSoul